# Towards Controllable Speech Synthesis in the Era of Large Language Models: A Survey

### v1

@@ IV-A2 Autoregressive (AR) Architectures
PromptStyle [229] VITS HiFi-GAN MelS 2023.05
StyleTTS 2 [89] Diffusion + GAN HifiGAN / iSTFTNet MelS 2023.06
VoiceBox [189] Flow Matching Diffusion HiFi-GAN MelS 2023.06
MegaTTS 2 [230] Diffusion + GAN HiFi-GAN MelS 2023.07
PromptTTS 2 [102] Diffusion Codec Decoder Token 2023.09
@@
DuIAN-E[232] CNN + RNN HiFi-GAN MelS 2023.09
PromptTTS++ [233] Transformer + Diffusion BigVGAN MelS 2023.09
SpeechFlow [190] Flow Matching Diffusion HiFi-GAN MelS 2023.10
P-Flow [188] Flow Matching HiFi-GAN MelS 2023.10
E3 TTS [184] Diffusion / Waveform→Unet 2023.11
HierSpeech++ [191] Hierarchical Conditional VAE BigVGAN MelS 2023.11
Audiobox [187] Flow Matching EnCodec MelS 2023.12
@@
NaturalSpeech 3 [87] Diffusion EnCodec Token 2024.04
@@
ControlSpeech [106] Transformer + Diffusion FACodec Decoder Token 2024.06
AST-LDM [234] Diffusion HiFi-GAN MelS 2024.06
SimpleSpeech [186] Transformer Diffusion SQ Decoder Token 2024.06
DiTTo-TTS [185] DiT BigVGAN Token 2024.06
E2 TTS [193] Flow Matching Transformer BigVGAN MelS 2024.06
MobileSpeech [235] ConFormer Decoder Vocos Token 2024.06
@@
ArtSpeech [236] RNN + CNN HiFI-GAN MelS+Energy+F0+TV 2024.07
CCSP [237] Diffusion Codec Decoder Token 2024.07
SimpleSpeech 2 [199] Flow-based Transformer Diffusion SQ Decoder Token 2024.08
E1 TTS [196] DiT BigVGAN Continuous Token 2024.09
VoiceGuider [238] Diffusion BigVGAN MelS 2024.09
StyleTTS-ZS [239] Diffusion + GAN HifiGAN / iSTFTNet Token 2024.09
NansyTTS [240] Transformer NANSY++ MelS 2024.09
@@
MS2KU-VTTS [242] Diffusion BigvGAN MelS 2024.10
MaskGCT [78] Masked Generative Transformers DAC + Vocos Token 2024.10
@@
Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). Timbre involves gender and age. MelS and LinS represent Mel Spectrogram and Linear Spectrogram respectively. TV represents the vocal tract variables proposed in  [236].
@@
GMVAE-Tacotron[130] CNN + RNN WaveRNN MelS 2018.12
@@
DurIAN[244] CNN + RNN MB-WaveRNN MelS 2019.09
@@
VALL-E [85] LLM EnCodec Token 2023.01
SpearTTS [203] LLM SoundStream Token 2023.02
VALL-E X [207] LLM EnCodec Token 2023.03
Make-a-voice [204] LLM BigVGAN Token 2023.05
TorToise [246] Transformer + DDPM Univnet MelS 2023.05
MegaTTS [91] LLM + GAN HiFi-GAN MelS 2023.06
SC VALL-E [205] LLM EnCodec Token 2023.07
Salle [247] LLM Codec Decoder Token 2023.08
UniAudio [248] LLM EnCodec Token 2023.10
ELLA-V [208] LLM EnCodec Token 2024.01
BaseTTS [216] LLM UnivNet Token 2024.02
ClaM-TTS [249] LLM BigVGAN MelS+Token 2024.04
RALL-E [209] LLM SoundStream Token 2024.05
ARDiT [220] Decoder-only Diffusion Transformer BigVGAN MelS 2024.06
VALL-E R [211] LLM Vocos Token 2024.06
VALL-E 2 [212] LLM Vocos Token 2024.06
Seed-TTS [217] LLM + Diffusion Transformer / Token 2024.06
VoiceCraft [93] LLM HiFi-GAN Token 2024.06
@@
XTTS [250] LLM + GAN HiFi-GAN MelS+Token 2024.06
CosyVoice [17] LLM + Conditional Flow Matching HiFi-GAN Token 2024.07
MELLE [219] LLM HiFi-GAN MelS 2024.07
Bailing TTS [218] LLM + Diffusion Transformer / Token 2024.08
VoxInstruct [103] LLM Vocos Token 2024.08
Emo-DPO [31] LLM HiFi-GAN Token 2024.09
@@
FireRedTTS [214] LLM + Conditional Flow Matching BigVGAN-v2 Token 2024.09
@@
CoFi-Speech [215] LLM BigVGAN Token 2024.09
Takin [206] LLM HiFi-Codec Token 2024.09
HALL-E [213] LLM EnCodec Token 2024.10
@@
Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). Timbre involves gender and age. MelS and LinS represent Mel Spectrogram and Linear Spectrogram respectively.
IV-B Control Strategies

### v2

@@ IV Controllable TTS
NaturalSpeech 2 [86] Diffusion RVQ-based Codec [86] Latent Feature 2023.04
PromptStyle [190] VITS + Flow HiFi-GAN MelS 2023.05
StyleTTS 2 [89] Flow-based Diffusion + GAN HifiGAN / iSTFTNet [191] MelS 2023.06
VoiceBox [192] Transformer + Flow HiFi-GAN MelS 2023.06
MegaTTS 2 [193] Decoder-only Transformer + GAN HiFi-GAN MelS 2023.07
PromptTTS 2 [102] Diffusion RVQ-based Codec Latent Feature 2023.09
@@
DurIAN-E[195] CNN + RNN HiFi-GAN MelS 2023.09
PromptTTS++ [104] Transformer + Diffusion BigVGAN [196] MelS 2023.09
SpeechFlow [197] Transformer + Flow HiFi-GAN MelS 2023.10
P-Flow [198] Transformer + Flow HiFi-GAN MelS 2023.10
E3 TTS [199] Diffusion Not required Waveform 2023.11
HierSpeech++ [200] Transformer + VAE + Flow BigVGAN MelS 2023.11
Audiobox [201] Transformer + Flow EnCodec [171] MelS 2023.12
@@
NaturalSpeech 3 [87] Transformer + Diffusion FACodec [87] Token 2024.04
@@
ControlSpeech [106] Transformer + Diffusion FACodec Token 2024.06
AST-LDM [203] Diffusion + VAE HiFi-GAN MelS 2024.06
SimpleSpeech [204] Transformer + Diffusion SQ Codec [204] Token 2024.06
DiTTo-TTS [205] DiT + VAE BigVGAN MelS 2024.06
E2 TTS [206] Transformer + Flow BigVGAN MelS 2024.06
MobileSpeech [207] Transformer Vocos [208] Token 2024.06
@@
ArtSpeech [210] RNN + CNN HiFI-GAN MelS + Energy + F0 2024.07
CCSP [211] Diffusion RVQ-based Codec [211] Token 2024.07
SimpleSpeech 2 [212] Flow-based DiT SQ Codec Token 2024.08
E1 TTS [213] DiT + Flow BigVGAN Token + MelS 2024.09
StyleTTS-ZS [214] Flow-based Diffusion + GAN Mel-based Decoder [214] MelS 2024.09
NansyTTS [215] Transformer NANSY++ [215] MelS 2024.09
@@
MS2KU-VTTS [217] Transformer BigVGAN MelS 2024.10
MaskGCT [78] Transformer + Flow Vocos Token 2024.10
EmoSphere++ [218] Transformer + Flow BigVGAN MelS 2024.11
EmoDubber [219] Transformer + Flow Flow-based Vocoder [219] MelS 2024.12
HED [220] Flow-based Diffusion Vocos MelS 2024.12
DiffStyleTTS [221] Transformer + Diffusion HiFi-GAN MelS 2025.01
DrawSpeech [222] Diffusion HiFi-GAN MelS 2025.01
ProEmo [223] Transformer HiFi-GAN MelS 2025.01
@@
Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). MelS: Mel Spectrogram. LinS: Linear Spectrogram.
IV-A Non-Autoregressive Architectures
In non-autoregressive TTS models, the model generates the entire output sequence $\mathbf{y}=(y_{1},y_{2},\dots,y_{T})$ at once, conditioned on the input sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$. The probability distribution for generating the sequence is:
$P(\mathbf{y}|\mathbf{x})=P(\mathbf{y}|\mathbf{x},\theta),$ (1)
Where $P(\mathbf{y}|\mathbf{x})$ is the likelihood of the output sequence $\mathbf{y}$ given the input $\mathbf{x}$, and $\theta$ represents the parameters of the model (e.g., weights). Since the output sequence is predicted simultaneously, the model learns to capture dependencies in a way that does not rely on previously generated outputs.
Transformer-based Approaches.  Advancements in controllable TTS technology highlight the integration of deep learning with audio processing, driven by Transformer-based architectures. Ren et al. [15] introduced FastSpeech, a feed-forward non-autoregressive Transformer model that significantly enhances TTS efficiency by reducing inference time and improving the stability issues found in autoregressive models like Tacotron 2. This model provides precise control over prosodic features through duration prediction, effectively tackling the one-to-many mapping challenge. FastSpeech 2[180] builds on this by integrating pitch and energy control, eliminating the need for the complex teacher-student distillation process, thus enhancing training efficiency and improving voice quality. Parallel Tacotron [84] further advances TTS by employing a variational autoencoder-based residual encoder, capturing intricate prosodic nuances. This approach, combined with iterative spectrogram loss, significantly enhances the naturalness and quality of synthesized speech. Additionally, FastPitch [77] incorporates direct pitch prediction into its architecture, enabling fully parallelized synthesis and precise pitch manipulation. This capability enhances expressiveness and retains the efficiency benefits established by FastSpeech. These innovations significantly contribute to the development of more interactive and natural AI-driven communication systems, underscoring the potential of integrating AI with human-centric disciplines to craft a future where technology and humanity coexist harmoniously.
VAE-based Approaches.  Recent advancements in controllable TTS systems are largely driven by the integration of VAE architectures, which enhance the flexibility and precision of speech modulation. Zhang et al.[18] pioneered the use of VAEs in end-to-end speech synthesis, creating disentangled latent representations that allow effective style control and transfer, especially in prosody and emotion management, outperforming the Global Style Token (GST) model [19] in style transfer tasks. Building on this, Hsu et al.[129] developed a hierarchical generative model with a conditional VAE framework and a Gaussian mixture model, enabling precise control over complex speech attributes such as environment and style, thus improving expressive speech synthesis through refined noise and speaker characteristic management. Liu et al.[188] further advanced the field with the CLONE model, a single-stage TTS system that resolves the one-to-many mapping issue and enhances high-frequency information reconstruction. By employing a conditional VAE with normalizing flows and a dual-path adversarial training mechanism with multi-band discriminators, CLONE achieves nuanced control over prosody and energy, demonstrating superior performance in both speech quality and prosody control compared to state-of-the-art models. These collective innovations highlight the adaptability of VAEs in managing complex speech generation tasks, marking significant progress toward more dynamic and versatile TTS technologies, with ongoing research promising even greater advancements.
Diffusion-based Approaches. The core concept of diffusion-based models is to generate target data by progressively removing noise. During the forward diffusion phase, noise is incrementally added to the original data to form a noise distribution. In the generation phase, a reverse denoising process is employed to gradually recover high-quality speech from the noise. Grad-StyleSpeech [189] introduces a hierarchical transformer encoder to create a representative noise prior distribution for speaker-adaptive settings using score-based diffusion models. NaturalSpeech 2 [86] uses a neural audio codec with residual vector quantizers to obtain quantized latent vectors, which are then generated using a diffusion model conditioned on text input. NaturalSpeech 3 [87] decomposes speech into distinct subspaces that represent different attributes and generates each subspace independently. DEX-TTS [209] improves DiT-based diffusion networks by applying overlapping patching and convolution-frequency patch embedding strategies. E3 TTS [199] models the temporal structure of the waveform through the diffusion process, eliminating the need for any intermediate representations, such as spectrogram features or alignment information.
Applying diffusion models to TTS requires a complex pipeline due to the need for precise temporal alignment between text and speech and the high fidelity required for audio data. This includes domain-specific modeling, such as phoneme and duration [86]. To address the issue of reduced naturalness caused by the addition of duration models, DiTTo-TTS [205] leverages the off-the-shelf pre-trained text and speech encoders without relying on speech domain-specific modeling by incorporating cross-attention mechanisms with the prediction of the total length of speech representations. Similarly, SimpleSpeech [204] proposes a speech codec model (SQ-Codec) based on scalar quantization and uses the sentence duration to control the generated speech length. Besides TTS, some text-to-audio models can also perform TTS and are worth referring to, such as AudioLDM [224], AudioLDM2 [225], Make-An-Audio [226], and CosyAudio [227].
Flow-based Approaches. Flow-based methods leverage invertible flow transformations [163, 228] to learn mappings from target speech features to simple distributions [179], typically standard Gaussian distributions. Due to their invertibility, this mechanism can directly sample from the simple distribution and generate high-fidelity speech in the reverse direction. Audiobox [201] and P-flow [198] employ non-autoregressive flow-matching models [228] for efficient and stable speech synthesis. VoiceBox [192] also employs flow-matching to generate speech, effectively casting the TTS task into a speech infilling task. SpeechFlow [197] is trained on 60k hours of untranscribed speech with flow matching and mask conditions and can be fine-tuned with task-specific data to match or surpass existing expert models. This highlights the potential of generative models as foundation models for speech applications. HierSpeech++ [200] proposes a hierarchical variational inference method. FlashSpeech [202] is built on a latent consistency model and applies a novel adversarial consistency training approach that can train from scratch without the need for a pre-trained diffusion model as the teacher, achieving speech generation in one or two steps.
Recently, E2 TTS [206] converts text input into a character sequence with filler tokens and trains a mel spectrogram generator based on an audio infilling task, achieving human-level naturalness. Inspired by E2 TTS, F5-TTS [229] refines the text representation with ConvNext v2 [230], facilitating easier alignment with speech. E1 TTS [213] further distills a diffusion-based TTS model into a one-step generator with distribution matching distillation [231, 232], reducing the number of network evaluations in sampling from diffusion models. SimpleSpeech 2 [212] introduces a flow-based scalar transformer diffusion model. The work also provides a theoretical analysis, showing that the inclusion of a small number of noisy labels in a large-scale dataset is equivalent to introducing classifier-free guidance during model optimization.
Other NAR Approaches. Other works leverage GAN-based or masked generative model-based methods for TTS generation. StyleTTS 2 [89] employs large pre-trained speech language models (SLMs) such as Wav2Vec 2.0 [165], HuBERT [166], and WavLM [233] as discriminators in combination with a novel differentiable duration modeling approach. This setup uses SLM representations to enhance the naturalness of the synthesized speech. MaskGCT [78] proposes masked generative transformers without requiring text-speech alignment supervision and phone-level duration prediction. The model employs a two-stage system trained using a mask-and-predict learning paradigm.
@@ IV-A Non-Autoregressive Architectures
GMVAE-Tacotron [129] VAE WaveRNN MelS 2018.12
@@
DurIAN [138] CNN + RNN Multi-band WaveRNN [138] MelS 2019.09
@@
VALL-E [85] Decoder-only Transformer EnCodec Token 2023.01
SpearTTS [237] Decoder-only Transformer SoundStream [170] Token 2023.02
VALL-E X [238] Decoder-only Transformer EnCodec Token 2023.03
Make-A-Voice [239] Encoder-decoder Transformer Unit-based Vocoder [239] Token 2023.05
TorToise [240] Decoder-only Transformer + Diffusion UnivNet [241] MelS 2023.05
MegaTTS [91] Decoder-only Transformer + GAN HiFi-GAN MelS 2023.06
SC VALL-E [242] Decoder-only Transformer EnCodec Token 2023.07
Salle [243] Decoder-only Transformer EnCodec Token 2023.08
UniAudio [244] Decoder-only Transformer EnCodec Token 2023.10
ELLA-V [245] Decoder-only Transformer EnCodec Token 2024.01
Base TTS [246] Decoder-only Transformer HiFi-GAN+BigVGAN Token 2024.02
CLaM-TTS [247] Encoder-decoder Transformer BigVGAN Token + MelS 2024.04
RALL-E [248] Decoder-only Transformer SoundStream Token 2024.05
ARDiT [249] Decoder-only DiT BigVGAN MelS 2024.06
VALL-E R [250] Decoder-only Transformer Vocos Token 2024.06
VALL-E 2 [251] Decoder-only Transformer Vocos Token 2024.06
Seed-TTS [252] Decoder-only Transformer + DiT Unknown Latent Feature 2024.06
VoiceCraft [93] Decoder-only Transformer HiFi-GAN Token 2024.06
@@
XTTS [253] Decoder-only Transformer HiFi-GAN-based Vocoder [253] Token + MelS 2024.06
CosyVoice [17] Decoder-only Transformer + Flow HiFi-GAN Token 2024.07
MELLE [254] Decoder-only Transformer HiFi-GAN MelS 2024.07
VoxInstruct [103] Decoder-only Transformer Vocos Token 2024.08
Emo-DPO [31] Decoder-only Transformer HiFi-GAN Token + MelS 2024.09
@@
FireRedTTS [255] Decoder-only Transformer + Flow BigVGAN Token + MelS 2024.09
@@
CoFi-Speech [256] Decoder-only Transformer BigVGAN Token + MelS 2024.09
Takin [257] Decoder-only Transformer + Flow HiFi-GAN Token + MelS 2024.09
HALL-E [258] Decoder-only Transformer EnCodec Token 2024.10
FishSpeech [259] Decoder-only Transformer Firefly-GAN [259] Token 2024.11
SLAM-Omni [260] Decoder-only Transformer HiFi-GAN Token + MelS 2024.12
IST-LM [261] Decoder-only Transformer HiFi-GAN Token + MelS 2024.12
KALL-E [262] Decoder-only Transformer WaveVAE [262] Latent Feature 2024.12
IDEA-TTS [263] Transformer Flow-based Vocoder [263] LinS + MelS 2024.12
FleSpeech [264] Flow-based DiT WaveGAN [141] Latent Feature 2025.01
Step-Audio [265] Decoder-only Transformer Flow-based Vocoder [265] Token 2025.02
Vevo [266] Decoder-only Transformer BigVGAN Token + MelS 2025.02
Spark-TTS [267] Decoder-only Transformer BiCodec [267] Token 2025.03
@@
Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). MelS: Mel Spectrogram. LinS: Linear Spectrogram.
IV-B Autoregressive Architectures
For an autoregressive model in TTS, the probability of the speech frame sequence $\mathbf{y}=(y_{1},y_{2},\dots,y_{T})$ given the input sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$ can be modeled as:
$P(\mathbf{y}|\mathbf{x})=\prod_{t=1}^{T}P(y_{t}|y_{<t},\mathbf{x}),$ (2)
where $y_{t}$ is the predicted output frame at time step $t$, $y_{<t}=(y_{1},\dots,y_{t-1})$ are the previous frames, and $\mathbf{x}$ is the input feature sequence. Each frame $y_{t}$ is predicted conditioned on all previous frames and the input sequence $\mathbf{x}$. Autoregressive models are powerful in TTS modeling but tend to be slower in generation time compared to non-autoregressive models, making them suitable for applications where quality is prioritized over real-time performance.
HMM-based Approaches.  In the realm of controllable TTS, advancements in HMM architectures have significantly enhanced the manipulation of speech elements such as emotion and prosody. Yamagishi et al. [70] pioneered this field by introducing style-dependent and style-mixed modeling, which allowed precise emulation of human-like emotional nuances and versatile synthesis across various styles by incorporating style as a contextual variable. Building on this foundation, Qin et al. [268] developed the “average emotion model,” which utilized maximum likelihood linear regression-based adaptation to modulate emotions like happiness and sadness even with limited data, thus advancing the emotional intelligence of synthetic speech systems.
Furthering expressive variability, Nose et al. [117] integrated subjective style intensities and a multiple-regression global variance model into HMMs, addressing over-smoothing and enabling nuanced emotional expressions. Lorenzo-Trueba et al. [72] expanded on these capabilities with CSMAPLR [72] adaptation, introducing “emotion transplantation” to transfer emotional states between speakers while preserving voice distinctiveness, enhancing personalized human-computer interaction. These innovations in HMM architectures have broadened the expressiveness and individuality in synthetic speech, augmenting technological interfaces and paving the way for future developments in adaptive, lifelike speech generation.
RNN-based Approaches. Controllable TTS technology has seen significant advancements through innovations in neural network architectures, particularly RNN-based architectures, enabling natural-sounding speech generation with adjustable emotion, prosody, and pitch. Prosody-Tacotron [234] is an extension of the original Tacotron model, designed to improve the prosody (rhythm, intonation, stress patterns) of synthesized speech. It builds on the Tacotron framework by introducing additional mechanisms to explicitly control prosodic features, which is a key challenge in TTS systems. Wang et al.[19] introduced global style tokens, using an unsupervised approach to encapsulate diverse speech styles into fixed tokens, thus enabling versatile style transfer within the Tacotron framework. Similarly, Stanton et al. [235] introduce Text-Predicted Global Style Tokens (TP-GST), enhancing Tacotron speech synthesis to predict expressive speaking styles directly from text without auxiliary inputs. Experiments show TP-GST generates more prosodic variation and achieves higher listener preference than baselines. Skerry-Ryan et al.[234] further advanced this by incorporating prosodic embeddings, providing control over timing and intonation, and significantly improving the replication of emotions in synthetic speech.
LLM-based Approaches. Inspired by the success of LLMs in natural language processing (NLP), recent studies have explored leveraging in-context learning for zero-shot TTS. As shown in Fig. 3, LLM-based approaches often take the target text or instructions and an optional reference speech clip as input, and utilize autoregressive modeling to generate speech tokens or features, which are then converted into the final waveform by a decoder.
[…]
Building on VALL-E, various improvements have been proposed. VALL-E X [238] extends VALL-E to multilingual scenarios, supporting zero-shot cross-lingual speech synthesis and speech-to-speech translation. ELLA-V [245] introduces a sequence order rearrangement step, enhancing local alignment between phoneme and acoustic modalities. RALL-E [248] incorporates prosody tokens as chain-of-thought prompting [270] to stabilize the generation of speech tokens. VALL-E R [250] improves phoneme-to-acoustic alignment and adopts codec-merging to boost decoding efficiency and reduce computational overhead. VALL-E 2 [251] introduces repetition-aware sampling and grouped code modeling for greater stability and faster inference. HALL-E [258] adopts a hierarchical post-training framework, effectively managing the trade-off between reducing the frame rate and producing high-quality speech.
Beyond the foundational improvements introduced by VALL-E and its extensions, further advancements have focused on enhancing speech alignment, quality, and robustness. SpearTTS [237] and Make-a-Voice [239] use semantic tokens to bridge the gap between text and acoustic features. FireRedTTS [255] further optimizes the tokenizer architecture to enhance speech quality. CoFi-Speech [256] generates speech in a coarse-to-fine manner via a multi-scale speech coding and generation approach, producing natural and intelligible speech. Similarly, BASE TTS [246] introduces discrete speech representations based on the WavLM [233] self-supervised model, focusing on phonemic and prosodic information. SeedTTS [252] also proposes a self-distillation method for speech decomposition and a reinforcement learning approach to enhance the robustness, speaker similarity, and controllability of generated speech.
Although models using discrete tokens as intermediate representations have achieved notable success in zero-shot TTS, they still face fidelity issues compared to continuous representations like Mel spectrograms [254, 249]. MELLE [254] optimizes the training objectives and sampling strategy, marking the first exploration of using continuous-valued tokens instead of discrete-valued tokens within the paradigm of autoregressive speech synthesis models. Similar to MELLE, ARDiT [249] encodes audio as a vector sequence in continuous space and autoregressively generates these sequences by a decoder-only transformer.
Additionally, some autoregressive methods enable speech editing through natural language instructions. VoiceCraft [93] introduces a decoder-only Transformer-based neural codec language model that utilizes causal masking and delayed stacking for token rearrangement. This approach allows for bidirectional context-aware speech editing and zero-shot TTS. VoiceCraft achieves precise control through text-guided modifications, such as insertion, deletion, and substitution, producing edits that are nearly indistinguishable from unmodified recordings in terms of naturalness. InstructSpeech [271] employs a multi-task LLM trained on triplet data (instruction, input, and output speech) with task embeddings and hierarchical adapters. It enables fine-grained control over both semantic attributes (content editing) and acoustic properties (emotion, speed) using natural language instructions. By leveraging multi-step reasoning, InstructSpeech facilitates free-form editing and efficiently adapts to new tasks.
IV-C Control Strategies